A computer randomly chooses a binary string of length 6. What is the probability that exactly two characters in the string turn out to be 1s?


1
Expert's answer
2021-11-23T18:02:56-0500

Let us use the formula for number of permutation with repetitions to find all strings of length 6 that exactly two characters in the strings turn out to be 1s. It follows that there are "P_6(2,4)=\frac{6!}{2!4!}=\frac{6\cdot 5}{2}=15" such strings. Taking into account that the total number of strings of length 6 is equal to "2^6=64," we conclude that the probability "p" that exactly two characters in the string turn out to be 1s is "p=\frac{15}{64}."